# Coherence-Assisted Entanglement Activation During Pure Dephasing Evolution

### v1

@@ Abstract
We investigate the system-environment entanglement activation during pure dephasing evolution, also called controlled unitary. A lower bound of generated entanglement is formulated with the initial coherence in the system and using relative entropy measure. For a qubit-environment case, this lower becomes a closed form by replacing conditional entropy with the overlap on the states in environment. Applying this result to the spin-boson model, we illustrate how the lower bound evolves as a function of time multiplied cutoff frequency, shown plots. The different behaviors of lower bounds with respect to Ohmicity, s=2 and s=3, corresponding to Markovian and non-Markovian evolution. In addition, we show that the decoherence in the system throughout the pure dephasing evolution represents another lower bound and upper bound of generated entanglement, indicating that larger entanglement generation leads more decoherence.
@@ 1 Introduction
Entanglement Horodecki et al. (2009) has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well. Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation. However, there are processes where entanglement can be generated spontaneously. Open systems dynamics (see e.g. Schlosshauer (2007); Breuer and Petruccione (2002) for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation.
Until recently Roszak and Cywiński (2015); Roszak (2018, 2020); Rzepkowski and Roszak (2020) this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence. However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea Zurek (2009) and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see Korbicz (2021) for a review), the environment has been recognized as a carrier of valuable information about the system. Therefore, studying system-environment entanglement is important not only for understanding decoherence processes, since entanglement in general increases decoherence efficiency, see e.g. Takahashi et al. (2022) for recent take on this question, but also to know how environment acquires system-related information.
One can illustrate system-environment entanglement generation on a paradigmatic example of the, so called, controlled unitary evolution, alternatively called also a measurement-type or a pure dephasing evolution, and defined by:
@@
in some preferred basis of the system $|s_{i}\rangle$, called the pointer basis  Zurek (1981, 1982, 2003). This basis is specified by the nature of the interaction, e.g. the dipole approximation for the electromagnetic interaction, leading to the spin-boson model Schlosshauer (2007); Breuer and Petruccione (2002). Whenever there is a superposition present in the initial state $|\psi_{S}\rangle=\sum_{i}a_{i}|s_{i}\rangle$, the resulting state $|\psi^{\prime}_{SE}\rangle$ is entangled:
@@
while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$. In the modern language of the resource theories Chitambar and Gour (2019), this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  Piani et al. (2011); Chuan et al. (2012); Streltsov et al. (2015); Killoran et al. (2016); Qiao et al. (2018); Regula et al. (2018). Let us remark here that an abstract coherence theory finds its natural application in open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
However easy the pure state interconversion (2) looks like, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement. Assuming that the initial state of the system is pure, Roszak et al.  Roszak and Cywiński (2015); Roszak (2018) recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement Roszak (2020); Rzepkowski and Roszak (2020). However, it is still unknown how to generalize their methods to mixed initial states of the system.
Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure Streltsov et al. (2015). However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case. In pure dephasing evolution, therefore, it is not clear how entanglement generation is quantified through the initial coherence or decoherence in the system.
In this work, we study entanglement generation for mixed states of the system. We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ Vedral et al. (1997); Vedral and Plenio (1998). Following the simple intuition (2) and its developments Piani et al. (2011); Chuan et al. (2012); Streltsov et al. (2015); Killoran et al. (2016); Qiao et al. (2018); Regula et al. (2018), the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
@@
Regarding the first lower bound, for two-dimensional central systems, i.e. for qubits, we derive a stronger result in given by a closed formula that links entanglement with the distinguishability of the environment states. In particular, this allows to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution. As an illustration, we apply our results to the spin-boson model (see e.g. Schlosshauer (2007) and the references therein). We study the dependence of the minimum generated entanglement on various model parameters e.g. the temperature and Ohmicity. Varying the latter through the well-known Markovian/non-Markovian transition, we see the traces of non-Markovianity in the dynamics of entanglement; cf. Rivas et al. (2010).
@@ 2 Preliminaries
$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
In order to evaluate the amount of $SE$ entanglement, we will use the relative entropy of entanglement $E_{r}$ Vedral et al. (1997); Vedral and Plenio (1998), defined as:
@@
where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$. This is a proper measure of entanglement Plenio and Virmani (2007), with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g. Vedral (2002). The minimization in definition is the reason why it is difficult to evaluate the value directly. We employ the following lemma to bypass this problem:
@@
Plenio et al. (2000) The relative entropy of entanglement is bounded from below by the negative conditional entropy, i.e. for any bipartite state $\rho_{AB}$ we have:
@@
We will also need elements of the quantum coherence theory; see e.g.  Baumgratz et al. (2014). Since quantum coherence is a basis dependent resource, we need to specify which basis is incoherent. In principle, one can develop a theory with any incoherent basis. However, since we study a concrete family of open dynamics, given by (3), the pointer basis of the central system $|i\rangle$ is the natural candidate here because if the system loses coherence due to the interaction, decoheres to this basis. We do not specify the incoherent basis for environment because the coherence in the environment does not matter in our analysis as we will show below. In the language of the coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology. We denote by $\mathcal{I}$ the set of all incoherent states. In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation respect to that resource. Various free operations for coherence are proposed for different physical motivations Chitambar and Gour (2016). The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}\in\mathcal{I}$ for all $n$ and any incoherent state $\rho\in\mathcal{I}$. One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) Baumgratz et al. (2014):
@@
Another ingredient that we will need is a measure of state distinguishability. Out of the possible measures Fuchs and van de Graaf (1999), we will use the state fidelity $B(\rho,\sigma)$, also called generalized overlap, defined as:
$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}=\textrm{Tr}\sqrt{\rho^{\frac{1}{2}}\sigma\rho^{\frac{1}{2}}}$ (9)
where $\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]$ is the probability that one obtains the same value $i$ from the measurement on $\rho$ and $\sigma$. Vanishing state fidelity is equivalent to a perfect distinguishability of the two states by POVM’s. This justifies the use of the state fidelity as a measure of distinguishability. This measure does not have as clear operational meaning as e.g. the quantum Chernoff bound Audenaert et al. (2007), related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (9) and it also provides both an upper and a lower bound for the quantum Chernoff information.
Discriminating states from an ensemble =$\mathcal{E}=\{p_{i},\rho_{i}\}$ is a more difficult task and its efficiency can be quantified by the, so called, accessible information $I_{acc}(\mathcal{E})$. It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ Cover and Thomas (2006) between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$. The formal expression is given by:
@@
Rahul and Ashwin (2007)[Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
@@ 3 Bounds on the system-environment entanglement generation
where $p_{i}=\langle i|\rho_{S}|i\rangle$ are the initial pointer probabilities and the conditional states of the environment $\rho_{i}(t)=U_{E}^{i}(t)\rho_{E}{U_{E}^{i}}(t)^{\dagger}$ encode, in a generally imperfect way, the information about the state of the central system. One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $i$, encoded by the environment in the course of the evolution. One may view this as an indirect measurement of the system through the environment, common in the macroscopic world; cf. the studies of objectivity in quantum mechanics in e.g. Zurek (2009); Korbicz (2021). The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ Cover and Thomas (2006). From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated. On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated Killoran et al. (2016). Indeed this is so as the following Lemma shows:
@@
$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m)=\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
@@
Finally, we will give an upper bound and another lower bound which we mentioned in the proof of lemma 3 on the generated entanglement. From a previous research Streltsov et al. (2015), an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
@@
Lemma 4 fits the intuition that larger system-environment entanglement leads to more decoherence quantified by $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$. In other words, decoherence in the system $D_{S}$ is lower bounded by generated entanglement.
@@ 4 Case study: The spin-boson model
In this Section, we illustrate our results on the spin-boson model, which is one of the most popular open systems models; see e.g. Schlosshauer (2007); Breuer and Petruccione (2007), and the references therein for an introduction. In its most used form, it describes a two-level central system, a qubit, interacting with a collection of bosonic modes (or harmonic oscillators), constituting the environment; for an example of higher dimensional central systems see e.g. Kiciński and Korbicz (2021). The dynamics of the model is given by the following Hamiltonian:
$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right),$ (26)
@@
where $\Lambda$ is the cutoff frequency and $s$ is the Ohmicity parameter. The value $s=3$ is a known threshold, distinguishing Markovian evolution for $s<3$ from non-Markovian for $s\geq 3$; see e.g. Addis et al. (2014). In Lampo et al. (2017) the analytical expression for $B$ was obtained for the case of the Ohmicity parameter $s>1$. It consists of two parts, vacuum, i.e. temperature-independent, and thermal:
@@
In Fig. 1, we keep fixed the ratio of the thermal to the cutoff energy, $T/\Lambda=1$, and plot (1) for different mixedness $\alpha$ and Ohmicity $s$ parameters. As expected from the (1), less mixed, i.e. more coherent, states with lower $\alpha$ lead to much more entanglement generation. More interestingly, there is a slight quantitative change of behavior between $s=2$ and $s=3$, corresponding to the transition between the Markovian and non-Markovian evolution respectively. While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless the value of $\alpha$ (depicted on the insert). This peak is more visible in Fig. 2, where we fix the mixedness $\alpha=1/2$ and vary the energy ratio $T/\Lambda$. The non-monotonic behavior of the entanglement bound for $s=3$ bears certain similarity to non-Markovianity measures, where a non-monotonic behavior of some marker function certifies information back-flow and hence the non-Markovianity of the evolution. Especially close seems the measure from Hall et al. (2014), where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.

### v2

# Coherence-Assisted Entanglement Activation During Open Evolution

@@ Abstract
The fact that in the course of an open evolution there is, in a vast majority of cases, a creation of system-environment entanglement out of the initial coherences, has been known for many years. However, how much entanglement can be created under general assumptions has not been analyzed yet. In this work, we investigate system-environment entanglement generation under a broad and important class of interactions known as pure dephasing interactions. Using completely general initials states, we derive a series of bounds on the amount of the created entanglement, connecting it to the initial system coherences as well as to the decohrence strength, making quantitatively precise the statement that larger entanglement leads to more decoherence. Interestingly, we obtain lower bounds, certifying minimal amount of entanglement that can be produced. As an illustration, we apply our results to the spin-boson model, where a qubit interacts with a bosonic bath, showing time evolution of the entanglement and the Markovian/non-Markovian crossover. Our results help to better understand and quantify the complicated correlations, produced during an open evolution, which will hopefully lead to better/new decoherence mitigation techniques.
@@ 1 Introduction
Entanglement [1] has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well. Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation. However, there are processes where entanglement can be generated or activated spontaneously. Open systems dynamics (see e.g. [2, 3] for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation out of initial system coherences.
Until recently [4, 5, 6, 7] this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence. However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea [8] and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see e.g. [9] for a review), the environment has been recognized as a carrier of valuable information about the system. Therefore, studying system-environment entanglement is important not only for understanding decoherence processes, since entanglement in general increases decoherence efficiency, see e.g. [10] for recent take on this question, but also to know how environment acquires system-related information.
One can illustrate the system-environment entanglement activation on a paradigmatic example of the, so called, pure dephasing evolution, also known as controlled-unitary evolution, alternatively called also a measurement-type evolution:
@@
in some preferred basis of the system $|s_{i}\rangle$, called the pointer basis  [11, 12, 13]. This basis is specified by the nature of the interaction, e.g. the dipole approximation for the electromagnetic interaction, leading to the spin-boson model [2, 3]. Whenever there is a coherence present in the initial state, $|\psi_{S}\rangle=\sum_{i}a_{i}|s_{i}\rangle$, the resulting state $|\psi^{\prime}_{SE}\rangle$ is obviously entangled:
@@
while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$. In the modern language of the resource theories [14], this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  [15, 16, 17, 18, 19, 20]. Let us remark here that an abstract coherence theory finds its natural application in open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
However easy the pure state interconversion (2) looks like, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement. Assuming that the initial state of the system is pure, Roszak et al.  [4, 5] recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement [6, 7]. However, it is still unknown how to generalize their methods to mixed initial states of the system.
Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure [17]. However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case. In pure dephasing evolution, therefore, it is not clear how entanglement generation is quantified through the initial coherence or decoherence in the system.
In this work, we study entanglement generation for mixed states of the system. We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ [21, 22]. Following the simple intuition (2) and its developments [15, 16, 17, 18, 19, 20], the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
@@
Regarding the first lower bound, for two-dimensional central systems, i.e. for qubits, we derive a stronger result in given by a closed formula that links entanglement with the distinguishability of the environment states. In particular, this allows to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution. As an illustration, we apply our results to the spin-boson model (see e.g. [2] and the references therein). We study the dependence of the minimum generated entanglement on various model parameters e.g. the temperature and Ohmicity. Varying the latter through the well-known Markovian/non-Markovian transition, we see the traces of non-Markovianity in the dynamics of entanglement; cf. [23].
@@ 2 Preliminaries
$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}$
$=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
In order to evaluate the amount of $SE$ entanglement, we will use the relative entropy of entanglement $E_{r}$ [21, 22], defined as:
@@
where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$. This is a proper measure of entanglement [24], with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g. [25]. The minimization in definition is the reason why it is difficult to evaluate the value directly. We employ the following lemma to bypass this problem:
@@
[26] The relative entropy of entanglement is bounded from below by the negative conditional entropy, i.e. for any bipartite state $\rho_{AB}$ we have:
@@
We will also need elements of the quantum coherence theory; see e.g.  [27]. Since quantum coherence is a basis dependent resource, we need to specify which basis is incoherent. In principle, one can develop a theory with any incoherent basis. However, since we study a concrete family of open dynamics, given by (3), the pointer basis of the central system $|i\rangle$ is the natural candidate here because if the system loses coherence due to the interaction, decoheres to this basis. We do not specify the incoherent basis for environment because the coherence in the environment does not matter in our analysis as we will show below. In the language of the coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology. We denote by $\mathcal{I}$ the set of all incoherent states. In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation respect to that resource. Various free operations for coherence are proposed for different physical motivations [28]. The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}\in\mathcal{I}$ for all $n$ and any incoherent state $\rho\in\mathcal{I}$. One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) [27]:
@@
Another ingredient that we will need is a measure of state distinguishability. Out of the possible measures [29], we will use the state fidelity $B(\rho,\sigma)$, also called generalized overlap, defined as:
$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}$
$=$ $\textrm{Tr}\sqrt{\rho^{\frac{1}{2}}\sigma\rho^{\frac{1}{2}}}$ (9)
where $\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]$ is the probability that one obtains the same value $i$ from the measurement on $\rho$ and $\sigma$. Vanishing state fidelity is equivalent to a perfect distinguishability of the two states by POVM’s. This justifies the use of the state fidelity as a measure of distinguishability. This measure does not have as clear operational meaning as e.g. the quantum Chernoff bound [30], related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (2) and it also provides both an upper and a lower bound for the quantum Chernoff information.
Discriminating states from an ensemble =$\mathcal{E}=\{p_{i},\rho_{i}\}$ is a more difficult task and its efficiency can be quantified by the, so called, accessible information $I_{acc}(\mathcal{E})$. It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ [31] between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$. The formal expression is given by:
@@
[32][Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
@@ 3 Bounds on the system-environment entanglement generation
where $p_{i}=\langle i|\rho_{S}|i\rangle$ are the initial pointer probabilities and the conditional states of the environment $\rho_{i}(t)=U_{E}^{i}(t)\rho_{E}{U_{E}^{i}}(t)^{\dagger}$ encode, in a generally imperfect way, the information about the state of the central system. One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $i$, encoded by the environment in the course of the evolution. One may view this as an indirect measurement of the system through the environment, common in the macroscopic world; cf. the studies of objectivity in quantum mechanics in e.g. [8, 9]. The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ [31]. From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated. On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated [18]. Indeed this is so as the following Lemma shows:
@@
$I_{acc}=$ $\max_{POVM:\{M_{m}\}}I(i:m)$
$=$ $\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
@@
Finally, we will give an upper bound and another lower bound which we mentioned in the proof of lemma 3 on the generated entanglement. From a previous research [17], an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
@@
Lemma 4 fits the intuition that larger system-environment entanglement leads to more decoherence as quantified by $D_{S}$. More precisely, we show that $D_{S}$ is lower bounded by the generated entanglement:
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})-H(i)\leq D_{S}.$ (26)
@@ 4 Case study: The spin-boson model
In this Section, we illustrate our results on the spin-boson model, which is one of the most popular open systems models; see e.g. [2, 3], and the references therein for an introduction. In its most used form, it describes a two-level central system, a qubit, interacting with a collection of bosonic modes (or harmonic oscillators), constituting the environment; for an example of higher dimensional central systems see e.g. [33]. The dynamics of the model is given by the following Hamiltonian:
$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right)$ (27)
@@
where $\Lambda$ is the cut-off frequency and $s$ is the Ohmicity parameter. The value $s=3$ is a known threshold, distinguishing Markovian evolution for $s<3$ from non-Markovian for $s\geq 3$; see e.g. [34]. In [35] the analytical expression for $B$ was obtained for the case of the Ohmicity parameter $s>1$. It consists of two parts, vacuum, i.e. temperature-independent, and thermal:
@@
In Fig. 1, we keep fixed the ratio of the thermal to the cutoff energy, $T/\Lambda=1$, and plot (1) for different mixedness $\alpha$ and Ohmicity $s$ parameters. As expected from the (1), less mixed, i.e. more coherent, states with lower $\alpha$ lead to much more entanglement generation. More interestingly, there is a slight quantitative change of behavior between $s=2$ and $s=3$, corresponding to the transition between the Markovian and non-Markovian evolution respectively. While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless the value of $\alpha$ (depicted on the insert). This peak is more visible in Fig. 2, where we fix the mixedness $\alpha=1/2$ and vary the energy ratio $T/\Lambda$. The non-monotonic behavior of the entanglement bound for $s=3$ bears certain similarity to non-Markovianity measures, where a non-monotonic behavior of some marker function certifies information back-flow and hence the non-Markovianity of the evolution. Especially close seems the measure from [36], where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.